# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ Abstract
Infinitely wide or deep neural networks (NNs) with independent and identically distributed (i.i.d.) parameters have been shown to be equivalent to Gaussian processes. Because of the favorable properties of Gaussian processes, this equivalence is commonly employed to analyze neural networks and has led to various breakthroughs over the years. However, neural networks and Gaussian processes are equivalent only in the limit; in the finite case there are currently no methods available to approximate a trained neural network with a Gaussian model with bounds on the approximation error. In this work, we present an algorithmic framework to approximate a neural network of finite width and depth, and with not necessarily i.i.d. parameters, with a mixture of Gaussian processes with error bounds on the approximation error. In particular, we consider the Wasserstein distance to quantify the closeness between probabilistic models and, by relying on tools from optimal transport and Gaussian processes, we iteratively approximate the output distribution of each layer of the neural network as a mixture of Gaussian processes. Crucially, for any NN and $\epsilon>0$ our approach is able to return a mixture of Gaussian processes that is $\epsilon$-close to the NN at a finite set of input points. Furthermore, we rely on the differentiability of the resulting error bound to show how our approach can be employed to tune the parameters of a NN to mimic the functional behavior of a given Gaussian process, e.g., for prior selection in the context of Bayesian inference. We empirically investigate the effectiveness of our results on both regression and classification problems with various neural network architectures. Our experiments highlight how our results can represent an important step towards understanding neural network predictions and formally quantifying their uncertainty.
@@ 1 Introduction
Deep neural networks have achieved state-of-the-art performance in a wide variety of tasks, ranging from image classification (Krizhevsky et al., 2012) to robotics and reinforcement learning (Mnih et al., 2013). In parallel with these empirical successes, there has been a significant effort in trying to understand the theoretical properties of neural networks (Goodfellow et al., 2016) and to guarantee their robustness (Szegedy et al., 2013). In this context, an important area of research is that of stochastic neural networks (SNNs), where some of the parameters of the neural network (weights and biases) are not fixed, but follow a distribution. SNNs include many machine learning models commonly used in practice, such as dropout neural networks (Gal and Ghahramani, 2016), Bayesian neural networks (Neal, 2012), neural networks with only a subset of stochastic layers (Favaro et al., 2023), and neural networks with randomized smoothing (Cohen et al., 2019). Among these, because of their convergence to Gaussian processes, particular theoretical attention has been given to infinite neural networks with independent and identically distributed (i.i.d.) parameters (Neal, 2012; Lee et al., 2017).
@@
In this paper, we propose an algorithmic framework to approximate the input-output distribution of an arbitrary SNN over a finite set of input points with a Gaussian Mixture Model (GMM), that is, a mixture of $M$ Gaussian distributions (McLachlan and Peel, 2000). Critically, the GMM approximation resulting from our approach comes with error bounds on its distance (in terms of the 2-Wasserstein distance [^1] Note that, as we will emphasize in Section 3, the choice of the $2$-Wasserstein distance to quantify the distance between a SNN and a GMM guarantees that a bound on the 2-Wasserstein distance also implies a bounds in how distant are their mean and variance. (Villani et al., 2009)) to the input-output distribution of the SNN. An illustrative example of our framework is shown in Figure 1(a), where, given a SNN trained on a 1D regression task (Figure 1(a)), our framework outputs a GMM approximation (Figure 1(b)) with error bounds on its closeness to the SNN (Figure 1(c)). Our approach is based on iteratively approximating the output distribution of each layer of the SNN with a mixture of Gaussian distributions and propagating this distribution through the next layer. In order to propagate a distribution through a non-linear activation function, we first approximate it with a discrete distribution, which we call a signature approximation. [^2] A discrete approximation of a continuous distribution is also called a codebook in the field of constructive quantization (Graf and Luschgy, 2007) or particle approximation in Bayesian statistics (Liu and Wang, 2016). The resulting discrete distribution can then be propagated exactly through a layer of the neural network (activation function and linear combination with weights and biases), which in the case of jointly Gaussian weights and biases, leads to a new Gaussian mixture distribution. To quantify the error between the SNN and the resulting GMM, we use techniques from optimal transport, probability theory, and interval arithmetic. In particular, for each of the approximation steps described above, we bound the error introduced in terms of the 2-Wasserstein distance. Then, we combine these bounds using interval arithmetic to bound the 2-Wasserstein distance between the SNN and the approximating GMM. Furthermore, by relying on the fact that GMMs can approximate any distribution arbitrarily well (Delon and Desolneux, 2020), we prove that by appropriately picking the parameters of the GMM, the bound on the approximation error converges uniformly to zero by increasing the size of the GMM. Additionally, we show that the resulting error bound is piecewise differentiable in the hyper-parameters of the neural network, which allows one to optimize the behavior of a NN to mimic that of a given GMM via gradient-based algorithms.
@@
We empirically validate our framework on various SNN architectures, including fully connected and convolutional layers, trained for both regression and classification tasks on various datasets including MNIST, CIFAR-10, and a selection of UCI datasets. Our experiments confirm that our approach can successfully approximate a SNN with a GMM with arbitrarily small error, albeit with increasing computational costs when the network’s depth increases. Furthermore, perhaps surprisingly, our results show that even a mixture with a relatively small number of components generally suffices to empirically approximate a SNN accurately. To showcase the importance of our results, we then consider two applications: (i) uncertainty quantification in SNNs, and (ii) prior selection for SNNs. In the former case we show how the GMM approximation resulting from our framework can be used to study and quantify the uncertainty of the SNN predictions in classification tasks on the MNIST and CIFAR-10 datasets. In the latter case, we consider prior selection for neural networks, which is arguably one of the most important problems in performing Bayesian inference with neural networks (Fortuin, 2022), and show that our framework allows one to precisely encode functional information in the prior of SNNs expressed as Gaussian processes (GPs), thereby enhancing SNNs’ posterior performance and outperforming state-of-the-art methods for prior selection of neural networks.
@@ 1.1 Related Works
The convergence of infinitely wide SNNs with i.i.d. parameters to GPs was first studied by Neal (2012) by relying on the central limit theorem. The corresponding GP kernel for the one hidden layer case was then analytically derived by Williams (1996). These results were later generalized to deep NNs (Hazan and Jaakkola, 2015; Lee et al., 2017; Matthews et al., 2018), convolutional layers (Garriga-Alonso et al., 2019; Novak et al., 2018; Garriga-Alonso and van der Wilk, 2021), and general non-linear activation functions (Hanin, 2023). However, these results only hold for infinitely wide or deep neural networks with i.i.d. parameters; in practice, NNs have finite size and depth and for trained NNs their parameters are generally not i.i.d.. To partially address these issues, recent works have started to investigate how these results apply in the finite case (Dyer and Gur-Ari, 2019; Antognini, 2019; Yaida, 2020; Bracale et al., 2021; Klukowski, 2022; Balasubramanian et al., 2024). In particular, Eldan et al. (2021) were the first to provide upper bounds on the rates at which a single layer SNN with a specific parameter distribution converges to the infinite width GP in terms of the Wasserstein distance. This result was later generalized to isotropic Gaussian weight distributions (Cammarota et al., 2023), deep Random NN (Basteri and Trevisan, 2024), and to other metrics, such as the total variation, sup norm, and Kolmogorov distances (Apollonio et al., 2023; Bordino et al., 2023; Favaro et al., 2023). However, to the best of our knowledge, all existing works are limited to untrained SNNs with i.i.d. weights and biases, which may be taken as priors in a Bayesian setting (Bishop and Nasrabadi, 2006) or may represent the initialization of gradient flows in an empirical risk minimization framework (Jacot et al., 2018). In contrast, critically, in our paper we allow for correlated and non-identical distributed weights and biases, thus also including trained posterior distributions of SNNs learned via Bayesian inference. In this context, Khan et al. (2019) showed that for a subset of Gaussian Bayesian inference techniques, the approximate posterior weight distributions are equivalent to the posterior distributions of GP regression models, implying a linearization of the approximate SNN posterior in function space (Immer et al., 2021). Instead, we approximate the SNN with guarantees in function space and our approach generalizes to any approximate inference technique.
@@
While the former set of works focuses on the convergence of SNNs to GPs, various works have considered the complementary problem of finding the distribution of the parameters of a SNN that mimic a given GP (Flam-Shepherd et al., 2017, 2018; Tran et al., 2022; Matsubara et al., 2021). This is motivated by the fact that GPs offer an excellent framework to encode functional prior knowledge; consequently, this problem has attracted significant interest in encoding functional prior knowledge into the prior of SNNs. In particular, closely related to our setting are the works of Flam-Shepherd et al. (2017) and Tran et al. (2022), which optimize the parametrization of the weight-space distribution of the SNN to minimize the KL divergence and 1-Wasserstein distance, respectively, with a desired GP prior. They then utilize the optimized weight prior to perform Bayesian inference, showing improved posterior performance. However, these methods lack formal guarantees on the closeness of the optimized SNN and the GP that the error bounds we derive in this paper provide.
@@ 2 Notation
For a vector ${{x}}\in{\mathbb{R}}^{n}$, we denote with $\|{{x}}\|$ the Euclidean norm on ${\mathbb{R}}^{n}$, and use ${{{x}}^{(i)}}$ to denote the $i$-th element of $x$. Similarly, for a matrix ${{M}}\in{\mathbb{R}}^{n\times m}$ we denote by $\|{{M}}\|$ the spectral (matrix) norm of ${{M}}$ and use ${{M}}^{(i,j)}$ to denote the $(i,j)$-th element of ${{M}}$. Further, ${{\bar{1}}}_{n}\in{\mathbb{R}}^{n}$ and ${{\bar{1}}}_{m\times n}\in{\mathbb{R}}^{m\times n}$, respectively, denote a vector and matrix of ones, and $I_{n}\in{\mathbb{R}}^{n\times n}$ is used to denote the identity matrix of size $n$. Given a linear transformation function (matrix) ${{T}}\in{\mathbb{R}}^{m\times n}$, the post image of a region ${\mathcal{X}}\subset{\mathbb{R}}^{n}$ under ${{T}}$ is defined as $\mathrm{Post}({\mathcal{X}},{{T}})=\{{{T}}{{x}}\mid{{x}}\in{\mathcal{X}}\}$. If ${\mathcal{X}}\subset{\mathbb{R}}^{n}$ is finite, we denote by $|{\mathcal{X}}|$ the cardinality of ${\mathcal{X}}$. Further, we define ${\mathbf{1}}_{\mathcal{X}}:{\mathbb{R}}^{n}\rightarrow\{0,1\}$ as the step function over region ${\mathcal{X}}\subset{\mathbb{R}}^{n}$, that is, ${\mathbf{1}}_{{\mathcal{X}}}({{x}})=\begin{cases}1&{{x}}\in{\mathcal{X}},\\ 0&\text{otherwise.}\end{cases}$
Given a measurable space $({\mathcal{X}},{\mathcal{B}}({\mathcal{X}}))$ with ${\mathcal{B}}({\mathcal{X}})$ being the $\sigma-$algebra, we denote with ${\mathcal{P}}({\mathcal{X}})$ the set of probability distributions on $({\mathcal{X}},{\mathcal{B}}({\mathcal{X}}))$. In this paper, for a metric space ${\mathcal{X}}$, ${\mathcal{B}}({\mathcal{X}})$ is assumed to be the Borel $\sigma$-algebra of ${\mathcal{X}}$. Considering two measurable spaces $({\mathcal{X}},{\mathcal{B}}({\mathcal{X}}))$ and $({\mathcal{Y}},{\mathcal{B}}({\mathcal{Y}}))$, a probability distribution $p\in{\mathcal{P}}({\mathcal{X}})$, and a measurable mapping $h:{\mathcal{X}}\rightarrow{\mathcal{Y}}$, we use $h\#p$ to denote the push-forward measure of $p$ by $h$, i.e., the measure on ${\mathcal{Y}}$ such that $\forall{\mathcal{A}}\in{\mathcal{Y}}$, $(h\#p)({\mathcal{A}}))=p(h^{-1}({\mathcal{A}}))$. For $N\in\mathbb{N}$, $\Pi_{N}=\{\pi\in{\mathbb{R}}^{N}_{\geq 0}:\sum^{N}_{i=1}{\pi^{{(i)}}}=1\}$ is the N-simplex. A discrete probability distribution $d\in{\mathcal{P}}({\mathcal{X}})$ is defined as $d=\sum_{i=1}^{N}{{{\pi}}^{(i)}}\delta_{{{c}}_{i}}$, where $\delta_{{c}}$ is the Dirac delta function centered at location ${{c}}\in{\mathcal{X}}$ and $\pi\in\Pi_{N}$. Lastly, the set of discrete probability distributions on ${\mathcal{X}}$ with at most $N$ locations is denoted as ${\mathcal{D}}_{N}({\mathcal{X}})\subset\mathcal{P}({\mathcal{X}})$.
@@ 3.1 Gaussian Processes and Gaussian Mixture Models
A Gaussian process (GP) $g$ is a stochastic process such that for any finite collection of input points ${\mathcal{X}}=\{{{x}}_{1},\ldots,{{x}}_{D}\}$ the joint distribution of $g({\mathcal{X}})\coloneqq\{g({{x}}_{1}),\ldots,g({{x}}_{D})\}$ follows a multivariate Guassian distribution with mean function ${m}$ and covariance function ${k}$, i.e., $g({\mathcal{X}})\sim\mathcal{N}({m}({\mathcal{X}})),{k}({\mathcal{X}},{\mathcal{X}}))$ (Adler and Taylor, 2009). A Gaussian Mixture Model (GMM) with $M\in\mathbb{N}$ components, is a set of $M$ GPs, also called components, averaged w.r.t. a probability vector $\pi\in\Pi_{M}$ (Tresp, 2000). Therefore, the probability distribution of a GMM follows a Gaussian mixture distribution.

### v2

@@ Abstract
Infinitely wide or deep neural networks (NNs) with independent and identically distributed (i.i.d.) parameters have been shown to be equivalent to Gaussian processes. Because of the favorable properties of Gaussian processes, this equivalence is commonly employed to analyze neural networks and has led to various breakthroughs over the years. However, neural networks and Gaussian processes are equivalent only in the limit; in the finite case there are currently no methods available to approximate a trained neural network with a Gaussian model with bounds on the approximation error. In this work, we present an algorithmic framework to approximate a neural network of finite width and depth, and with not necessarily i.i.d. parameters, with a mixture of Gaussian processes with bounds on the approximation error. In particular, we consider the Wasserstein distance to quantify the closeness between probabilistic models and, by relying on tools from optimal transport and Gaussian processes, we iteratively approximate the output distribution of each layer of the neural network as a mixture of Gaussian processes. Crucially, for any NN and $\epsilon>0$ our approach is able to return a mixture of Gaussian processes that is $\epsilon$-close to the NN at a finite set of input points. Furthermore, we rely on the differentiability of the resulting error bound to show how our approach can be employed to tune the parameters of a NN to mimic the functional behavior of a given Gaussian process, e.g., for prior selection in the context of Bayesian inference. We empirically investigate the effectiveness of our results on both regression and classification problems with various neural network architectures. Our experiments highlight how our results can represent an important step towards understanding neural network predictions and formally quantifying their uncertainty.
@@ 1 Introduction
Deep neural networks have achieved state-of-the-art performance in a wide variety of tasks, ranging from image classification (Krizhevsky et al., 2012) to robotics and reinforcement learning (Mnih et al., 2013). In parallel with these empirical successes, there has been a significant effort in trying to understand the theoretical properties of neural networks (Goodfellow et al., 2016) and to guarantee their robustness (Szegedy et al., 2013). In this context, an important area of research is that of stochastic neural networks (SNNs), where some of the parameters of the neural network (weights and biases) are not fixed, but follow a distribution. SNNs include many machine learning models commonly used in practice, such as dropout neural networks (Gal and Ghahramani, 2016), Bayesian neural networks (Neal, 2012), neural networks with only a subset of stochastic layers (Favaro et al., 2025), and neural networks with randomized smoothing (Cohen et al., 2019). Among these, because of their convergence to Gaussian processes, particular theoretical attention has been given to infinite neural networks with independent and identically distributed (i.i.d.) parameters (Neal, 2012; Lee et al., 2018).
@@
In this paper, we propose an algorithmic framework to approximate the input-output distribution of an arbitrary SNN over a finite set of input points with a Gaussian Mixture Model (GMM), that is, a mixture of $M$ Gaussian distributions (McLachlan and Peel, 2000). Critically, the GMM approximation resulting from our approach comes with error bounds on its distance (in terms of the 2-Wasserstein distance [^1] Note that, as we will emphasize in Section 3, the choice of the $2$-Wasserstein distance to quantify the distance between a SNN and a GMM guarantees that a bound on the 2-Wasserstein distance also implies a bounds in how distant are their mean and variance. (Villani et al., 2009)) to the input-output distribution of the SNN. An illustrative example of our framework is shown in Figure 1(a), where, given a SNN trained on a 1D regression task (Figure 1(a)), our framework outputs a GMM approximation (Figure 1(b)) with error bounds on its closeness to the SNN (Figure 1(c)). Our approach is based on iteratively approximating the output distribution of each layer of the SNN with a mixture of Gaussian distributions and propagating this distribution through the next layer. In order to propagate a distribution through a non-linear activation function, we first approximate it with a discrete distribution, which we call a signature approximation. [^2] A discrete approximation of a continuous distribution is also called a codebook in the field of constructive quantization (Graf and Luschgy, 2000) or particle approximation in Bayesian statistics (Liu and Wang, 2016). The resulting discrete distribution can then be propagated exactly through a layer of the neural network (activation function and affine combination with weights and biases), which in the case of jointly Gaussian weights and biases, leads to a new Gaussian mixture distribution. To quantify the error between the SNN and the resulting GMM, we use techniques from optimal transport, probability theory, and interval arithmetic. In particular, for each of the approximation steps described above, we bound the error introduced in terms of the 2-Wasserstein distance. Then, we combine these bounds using interval arithmetic to bound the 2-Wasserstein distance between the SNN and the approximating GMM. Furthermore, by relying on the fact that GMMs can approximate any distribution arbitrarily well (Delon and Desolneux, 2020), we prove that by appropriately picking the parameters of the GMM, the bound on the approximation error converges uniformly to zero by increasing the size of the GMM. Additionally, we show that the resulting error bound is piecewise differentiable in the hyper-parameters of the neural network, which allows one to optimize the behavior of a NN to mimic that of a given GMM via gradient-based algorithms.
@@
We empirically validate our framework on various SNN architectures, including fully-connected and convolutional layers, trained for both regression and classification tasks on various datasets including MNIST, CIFAR-10, and a selection of UCI datasets. Our experiments confirm that our approach can successfully approximate a SNN with a GMM with arbitrarily small error, albeit with increasing computational costs when the network’s depth increases. Furthermore, perhaps surprisingly, our results show that even a mixture with a relatively small number of components generally suffices to empirically approximate a SNN accurately. To showcase the importance of our results, we then consider two applications: (i) uncertainty quantification in SNNs, and (ii) prior selection for SNNs. In the former case we show how the GMM approximation resulting from our framework can be used to study and quantify the uncertainty of the SNN predictions in classification tasks on the MNIST and CIFAR-10 datasets. In the latter case, we consider prior selection for neural networks, which is arguably one of the most important problems in performing Bayesian inference with neural networks (Fortuin, 2022), and show that our framework allows one to precisely encode functional information in the prior of SNNs expressed as Gaussian processes (GPs), thereby enhancing SNNs’ posterior performance and outperforming state-of-the-art methods for prior selection of neural networks.
@@ 1.1 Related Works
The convergence of infinitely wide SNNs with i.i.d. parameters to GPs was first studied by Neal (2012) by relying on the central limit theorem. The corresponding GP kernel for the one hidden layer case was then analytically derived by Williams (1996). These results were later generalized to deep NNs (Hazan and Jaakkola, 2015; Lee et al., 2018; Matthews et al., 2018), convolutional layers (Garriga-Alonso et al., 2019; Novak et al., 2018; Garriga-Alonso and van der Wilk, 2021), and general non-linear activation functions (Hanin, 2023). However, these results only hold for infinitely wide or deep neural networks with i.i.d. parameters; in practice, NNs have finite size and depth and for trained NNs their parameters are generally not i.i.d.. To partially address these issues, recent works have started to investigate how these results apply in the finite case (Dyer and Gur-Ari, 2020; Antognini, 2019; Yaida, 2020; Bracale et al., 2021; Klukowski, 2022; Balasubramanian et al., 2024). In particular, Eldan et al. (2021) were the first to provide upper bounds on the rates at which a single layer SNN with a specific parameter distribution converges to the infinite width GP in terms of the Wasserstein distance. This result was later generalized to isotropic Gaussian parameter distributions (Cammarota et al., 2023), deep Random NN (Basteri and Trevisan, 2024), and to other metrics, such as the total variation, sup norm, and Kolmogorov distances (Apollonio et al., 2025; Bordino et al., 2023; Favaro et al., 2025). However, to the best of our knowledge, all existing works are limited to untrained SNNs with i.i.d. weights and biases, which may be taken as priors in a Bayesian setting (Bishop and Nasrabadi, 2006) or may represent the initialization of gradient flows in an empirical risk minimization framework (Jacot et al., 2018). In contrast, critically, in our paper we allow for correlated and non-identical distributed weights and biases, thus also including trained posterior distributions of SNNs learned via Variational Inference. In this context, Khan et al. (2019) showed that for a subset of Gaussian Bayesian inference techniques, the approximate posterior parameter distributions are equivalent to the posterior distributions of GP regression models, implying a linearization of the approximate SNN posterior in function space (Immer et al., 2021). Instead, we approximate the SNN with guarantees in function space and our approach generalizes to any approximate inference technique.
@@
While the former set of works focuses on the convergence of SNNs to GPs, various works have considered the complementary problem of finding the distribution of the parameters of a SNN that mimic a given GP (Flam-Shepherd et al., 2017, 2018; Tran et al., 2022; Matsubara et al., 2021). This is motivated by the fact that GPs offer an excellent framework to encode functional prior knowledge; consequently, this problem has attracted significant interest in encoding functional prior knowledge into the prior of SNNs. In particular, closely related to our setting are the works of Flam-Shepherd et al. (2017) and Tran et al. (2022), which optimize the parametrization of the parameter-space distribution of the SNN to minimize the KL divergence and 1-Wasserstein distance, respectively, with a desired GP prior. They then utilize the optimized parameter prior to perform Bayesian inference, showing improved posterior performance. However, these methods lack formal guarantees on the closeness of the optimized SNN and the GP that the error bounds we derive in this paper provide.
@@ 2 Notation
For a vector ${{x}}\in{\mathbb{R}}^{n}$, we denote with $\|{{x}}\|$ the Euclidean norm on ${\mathbb{R}}^{n}$, and use ${{x}}^{(i)}$} to denote the $i$-th element of $x$. Similarly, for a matrix ${{M}}\in{\mathbb{R}}^{n\times m}$ we denote by $\|{{M}}\|$ the spectral (matrix) norm of M{{M}} and use ${{M}}^{(i,j)}$ to denote the $(i,j)$-th element of M{{M}}. In addition, $\mathrm{diag}({{x}})$ denotes a matrix with the elements of x{{x}} on its diagonal. Given a linear transformation function (matrix) ${{T}}\in{\mathbb{R}}^{m\times n}$, the post image of a region ${\mathcal{X}}\subset{\mathbb{R}}^{n}$ under T{{T}} is defined as $\mathrm{Post}({\mathcal{X}},{{T}})=\{{{T}}{{x}}\mid{{x}}\in{\mathcal{X}}\}$. If $\mathcal{X}$} is finite, we denote by $|{\mathcal{X}}|$ the cardinality of $\mathcal{X}$}.
Given a measurable space $({\mathcal{X}},{\mathcal{B}}({\mathcal{X}}))$ with ${\mathcal{B}}({\mathcal{X}})$ being the $\sigma$-algebra, we denote with ${\mathcal{P}}({\mathcal{X}})$ the set of probability distributions on $({\mathcal{X}},{\mathcal{B}}({\mathcal{X}}))$. In this paper, for a metric space $\mathcal{X}$}, ${\mathcal{B}}({\mathcal{X}})$ is assumed to be the Borel $\sigma$-algebra of $\mathcal{X}$}. Considering two measurable spaces $({\mathcal{X}},{\mathcal{B}}({\mathcal{X}}))$ and $({\mathcal{Y}},{\mathcal{B}}({\mathcal{Y}}))$, a probability distribution $p\in{\mathcal{P}}({\mathcal{X}})$, and a measurable mapping $h:{\mathcal{X}}\rightarrow{\mathcal{Y}}$, we use h#ph\#p to denote the pushforward measure of $p$ by $h$, i.e., the measure on $\mathcal{Y}$} such that $\forall{\mathcal{A}}\in{\mathcal{B}}({\mathcal{Y}})$, $(h\#p)({\mathcal{A}}))=p(h^{-1}({\mathcal{A}}))$. For $N\in\mathbb{N}$, $\Pi_{N}=\{\pi\in{\mathbb{R}}^{N}_{\geq 0}:\sum^{N}_{i=1}{\pi^{{(i)}}}=1\}$ is the N-simplex. A discrete probability distribution $d\in{\mathcal{P}}({\mathcal{X}})$ is defined as $d=\sum_{i=1}^{N}{{{\pi}}^{(i)}}\delta_{{{c}}_{i}}$, where $\delta_{{c}}$ is the Dirac delta function centered at location ${{c}}\in{\mathcal{X}}$ and $\pi\in\Pi_{N}$. Lastly, the set of discrete probability distributions on $\mathcal{X}$} with at most $N$ locations is denoted as ${\mathcal{D}}_{N}({\mathcal{X}})\subset\mathcal{P}({\mathcal{X}})$.
@@ 3.1 Gaussian Processes and Gaussian Mixture Models
A Gaussian process (GP) $g$ is a stochastic process such that for any finite collection of input points ${\mathcal{X}}=\{{{x}}_{1},\ldots,{{x}}_{D}\}$ the joint distribution of $g({\mathcal{X}})\coloneqq\{g({{x}}_{1}),\ldots,g({{x}}_{D})\}$ follows a multivariate Guassian distribution with mean function m{m} and covariance function k{k}, i.e., $g({\mathcal{X}})\sim\mathcal{N}({m}({\mathcal{X}})),{k}({\mathcal{X}},{\mathcal{X}}))$ (Adler and Taylor, 2009). A Gaussian Mixture Model (GMM) with $M\in\mathbb{N}$ components, is a set of $M$ GPs, also called components, averaged w.r.t. a probability vector $\pi\in\Pi_{M}$ (Tresp, 2000). Therefore, the probability distribution of a GMM follows a Gaussian mixture distribution.